A survey of college sophmores showed that 5/6 of the males were taking a math class and 3/4 of the females were taking a math class. One third of the males were enrolled in calcus, and 1/5 of the females were enrolled in calculus. If just as many males as females were surveyed, then what fraction of the surveyed students taking math were enrolled in calculus? Rework this problem assuming that the numbers of females in the survey was twice the number of males.
Answer by stanbon(75887)

Let "m"be the number of males; Let "f" be the number of females.
Rework this problem assuming that the number of females in the survey was twice the number of males.
Therefore f = 2m
-------------------------------------
A survey of college sophomores showed that 5/6 of the males were taking a math class and 3/4 of the females were taking a math class.
Number taking a math class: (5/6)m+(3/4)(2m) = (5/6 + 3/2)m = (7/3)m
--------------------------
One third of the males were enrolled in calculus, and 1/5 of the females were enrolled in calculus.
Number taking calculus: (1/3)m+(1/5)(2m) = (1/3 + 2/5)m = (11/15)m
-------------
Fraction taking calculus:
[(11/15)m]/[(7/3)m] = 33/105
=================================
Cheers,
Stan H.

If just as many males as females were surveyed, then what fraction of the surveyed students taking math were enrolled in calculus?
Rework this problem assuming that the numbers of females in the survey was twice the number of males.